# Phase Transition for Budgeted Multi-Agent Synergy

### v1

@@ Abstract
Multi-agent systems can improve reliability, yet under a fixed inference budget they often help, saturate, or even collapse. We develop a deliberately minimal and calibratable theory that predicts these regimes from three binding constraints of modern agent stacks: finite context windows, lossy inter-agent communication, and shared failures among similar agents. Each leaf agent is summarized by a compute–performance scaling exponent $\beta$; communication is captured by a message-length fidelity curve $\gamma(m)$; dependence is captured by an effective shared-error correlation $\rho$; and a context window $W$ imposes hard fan-in limits that make hierarchy necessary. For binary success/failure tasks with majority aggregation, we prove a sharp phase transition for deep $b$-ary trees with correlated inputs and lossy communication: a single scalar $\alpha_{\rho}$ (combining $\gamma(m)$, $\rho$, and fan-in $b$) determines whether weak signal is amplified to a nontrivial fixed point or washed out to chance. In the amplifying regime, we derive an organization exponent $s$ and show that budgeted synergy, i.e., outperforming the best single agent under the same total budget, occurs exactly when $s>\beta$, yielding closed-form compute allocation rules and explicit budget thresholds. We further characterize saturation via a mixing depth and provide a conservative clipped predictor that remains accurate across growth and saturation. A continuous-performance warm-up gives closed-form risks for star, chain, and tree organizations, making correlation- and communication-induced floors explicit and exposing the core design trade-offs in a smooth setting. Finally, we validate the predicted phase boundaries in controlled synthetic simulations and show how the same mechanisms explain the dominant bottlenecks reported in recent large-scale matched-budget studies of LLM agent-system scaling, including context saturation, subcritical error cascades, and diminishing returns at strong baselines.
@@ 1 Introduction
Multi-agent systems are often sold as a reliability primitive: if a single agent is fallible, run many agents, let them interact, and aggregate their outputs. Under a fixed inference budget, however, multi-agent coordination is not reliably beneficial. The same additional agents that sometimes improve performance can also yield saturation or outright degradation. This paper develops a predictive theory of this brittleness: when multi-agent organization produces budgeted synergy (i.e., outperforming any single agent under the same budget) and when scaling out must fail.
@@
The LLM era has made the phenomenon impossible to ignore. It is now routine to instantiate multiple LLM-based agents, assign roles, and coordinate them through discussion, critique, or voting in “agent societies” and orchestration toolkits (Brown et al., 2020; Yao et al., 2023; Wu et al., 2023; Li et al., 2023; Chen et al., 2024; Hong et al., 2024). The underlying intuition is classical: ensembling can reduce error (Breiman, 1996; Freund and Schapire, 1997; Dietterich, 2000), and interaction protocols such as debate or amplification aim to turn many weak judgments into a stronger one (Irving et al., 2018; Christiano et al., 2018; Du et al., 2024). At the same time, controlled matched-budget evaluations increasingly report strong task- and topology-dependence, including regimes where multi-agent variants degrade once coordination overhead and information loss dominate (Kim et al., 2025).
@@
Our starting point is that these failures are not implementation accidents but structural consequences of three constraints that repeatedly bind in modern agent stacks. First, agent errors are not independent: many instances derived from the same base model and prompt share failure modes, producing “groupthink” rather than error cancellation. Second, coordination is communication-limited: as systems scale, agents must compress reasoning into bounded-length messages, and this loss compounds across layers. Third, coordination is context-limited: a central aggregator cannot read arbitrarily many messages under a finite context window, so naive scale-out saturates structurally. Recent work on large-population collaboration and topology learning highlights the same bottleneck: token and context limits are often the binding resource (Qian et al., 2025; Zhang et al., 2024).
@@
Today, agent-system design still largely proceeds by heuristics: choose a topology, message format, and number of agents, then iterate. In spirit, this resembles the hand-crafted feature engineering era of machine learning: powerful components exist, but system-level behavior depends on ad hoc choices and expensive trial-and-error. Our goal is to push multi-agent design toward an “agentic intelligence from first principles” style: make the dominant constraints explicit, explain why scaling out works or fails, and derive sharp, quantitative regime boundaries—in particular, a phase transition for budgeted multi-agent synergy.
@@
Within this framework we analyze three canonical organizations: star (central aggregation), chain (sequential handoff), and hierarchical tree. Our main results characterize when hierarchy is viable and when it is provably counterproductive. For binary success/failure tasks with majority aggregation, we prove a sharp amplification–collapse transition for deep trees: a scalar $\alpha_{\rho}$ (combining communication fidelity, correlation, and fan-in) determines whether weak signal amplifies across layers or is washed out to chance, in the spirit of reconstruction thresholds in noisy broadcasting on trees (Evans et al., 2000; Mossel, 2001; Kesten and Stigum, 1966). In the amplifying regime, we derive an organization exponent $s$ governing growth with the number of leaves and show that budgeted synergy in the growth range occurs exactly when $s>\beta$, yielding closed-form compute allocation and explicit budget thresholds. We also characterize finite-depth saturation via a nontrivial fixed point and mixing-depth guarantees. Finally, we make precise why topology becomes decisive only because constraints bind: absent context limits, central aggregation is information-dominant (data processing inequality) (Cover and Thomas, 2006), and hierarchy is valuable only insofar as it bypasses fan-in limits without crossing into the collapsing regime.
@@
This is a theory paper: our aim is to delineate regimes of behavior and identify the mechanisms that govern when multi-agent systems help, saturate, or collapse under realistic constraints. Large-scale matched-budget evaluations with real LLM agent stacks are expensive and can be confounded by prompt engineering and rapidly changing model/tool ecosystems. Accordingly, we validate the sharp regime boundaries via controlled synthetic simulations (Section 7.2) that probe the theory under its own assumptions, and we use external controlled evidence (Kim et al., 2025) to show that the bottlenecks isolated by the theory (context saturation, subcritical collapse, and diminishing returns) are salient in modern deployments.
@@
Section 2 defines the framework, measurable parameters, and the budget and context constraints. Section 3 develops the continuous warm-up. Section 4 presents the binary theory, including the phase transition, small-signal amplification, and saturation. Section 5 compares topologies under context constraints and derives budget thresholds. Section 6 turns the theory into design diagnostics. Section 7 provides empirical touchpoints via external controlled evidence and synthetic sanity checks. We place detailed proofs and optional calibration/evaluation templates in the appendices.
@@ 2.1 Tasks and performance metrics
We study two task primitives that cover many MAS evaluations and isolate the coordination bottlenecks above. Binary success/failure captures accuracy-like outcomes (pass rates, accept/reject decisions). Continuous performance on $[0,1]$ captures graded scoring or calibrated estimates and provides a smooth warm-up where many expressions admit closed forms. The continuous setting is not a substitute for the binary theory; it makes correlation and communication floors explicit without the nonlinearity introduced by majority vote.
@@
$\mu\;:=\;\mathbb{E}[\widehat{Y}Y]\in[-1,1],\qquad\Pr(\widehat{Y}=Y)=\frac{1+\mu}{2}.$
The bias $\mu$ is convenient for multi-layer analysis because it composes cleanly under simple channel models and aggregation maps.
@@
$v\;:=\;\mathbb{E}\big[(\widehat{Y}-Y)^{2}\big],$
and occasionally report a bounded performance score such as $\mathrm{Perf}=1/(1+v)$, which is monotone in $v$. The continuous analysis yields explicit floors; the binary analysis additionally introduces a sharp amplification–collapse transition.
@@ 2.2 Single-agent capability and scaling
Our central comparison is fixed-budget: the same total budget $B$ can be spent on fewer stronger agents (scale up) or more weaker agents (scale out). We model per-leaf compute by a single knob $x\geq 0$. In an LLM agent, $x$ can represent a token budget for reasoning, the number of samples or self-consistency votes, the number of tool calls, or any additive resource that trades off against parallelism.
We summarize the effect of compute via a single-agent scaling law, motivated by empirical power-law relations between performance and compute observed for large language models and related systems (Kaplan et al., 2020; Hoffmann et al., 2022). For binary tasks, a leaf with compute $x$ produces a vote $\widehat{Y}$ with bias
@@
where $g(\cdot)$ is increasing and typically exhibits diminishing returns. For continuous tasks, a leaf produces an estimate $X$ with conditional variance $v_{0}(x)$, where $v_{0}(\cdot)$ decreases with $x$.
@@
with constants $k,c>0$ and a single-agent scaling exponent $\beta>0$. The exponent $\beta$ is not a universal constant. It depends on the model, prompting, tools, and the task family, and it is intended to be measured.
@@ 2.3 Communication as a controllable lossy channel
For binary tasks and majority-style protocols, it is natural to treat each child as intending to transmit a single bit of information (its vote or decision), encoded into a message of length $m$. We model the receiver’s decoded bit as passing through a binary symmetric channel (BSC) with reliability $\gamma(m)\in(0,1]$: if a child has bias $\mu$ relative to $Y$, then after transmission the received bit has bias $\gamma(m)\mu$. Equivalently, $\gamma(m)=1-2\epsilon(m)$ where $\epsilon(m)$ is an effective flip probability.
@@
$\widetilde{X}=X+\eta,\qquad\mathbb{E}[\eta]=0,\quad\mathrm{Var}(\eta)=\sigma_{c}^{2}(m),$
@@ 2.5 Budget and context constraints
$\text{fan-in}\leq\Big\lfloor\frac{W}{m}\Big\rfloor.$
@@ 2.6 Topologies and protocols
[…]
Figure 1: Canonical organizations under a context window constraint $W$. A star requires the root to read all messages, so feasibility enforces $Nm\leq W$. A hierarchical tree enforces the constraint locally ($bm\leq W$), enabling many more leaves at the cost of multi-hop communication and repeated aggregation. Chains avoid fan-in bottlenecks but compound communication loss.
@@
It is also useful to view more general feed-forward organizations through these motifs. With a finite context window, every node can only aggregate a bounded number of incoming messages, i.e., each local computation is an in-star with fan-in at most $\lfloor W/m\rfloor$. Globally, any connected directed acyclic communication graph contains a spanning arborescence that routes leaf information to the output; star and chain are the two extreme cases (depth $1$ versus fan-in $1$), and bounded-fan-in trees are the minimal structure that scales beyond the star’s context cap. Additional edges or multi-round interaction are valuable mainly when they effectively change the measurable parameters, for example by improving effective fidelity through redundancy or by reducing effective correlation through independent verification, rather than by rerouting the same information along multiple dependent paths.
@@ 3.1 Model assumptions for the continuous setting
Let $Y\in[0,1]$ denote the latent truth and let each leaf output an estimate $X\in[0,1]$. To keep the warm-up focused on organization rather than estimator bias, we assume leaf estimates are conditionally unbiased:
$\mathbb{E}[X\mid Y]=Y.$
We write the leaf residual as E:=XYE:=X-Y, so $\mathbb{E}[E\mid Y]=0$. Leaf quality is summarized by the conditional MSE,
$v_{0}(x):=\mathbb{E}\big[(X-Y)^{2}\big]=\mathbb{E}\big[E^{2}\big],$
which depends on per-leaf compute $x$. In the scaling regime of interest we model
@@
with constants $c>0$ and $\beta>0$, as introduced in Section 2.2.
Communication distortion.
An estimate sent across an edge as a message of length $m$ is received with additive distortion:
where $\sigma_{c}^{2}(m)$ decreases with $m$. We treat the channel noise variables $\eta$ as independent across edges and independent of $Y$. This is a compact abstraction of imperfect summarization, constrained prompts, and the general fact that receivers do not reconstruct sender information perfectly.
Shared errors as residual correlation.
Within any aggregation group, child residuals are not assumed independent. Instead, we use an equal-correlation model: for children $i\neq j$ with residuals $E_{i},E_{j}$,
@@
where $\rho\in[0,1)$ is the correlation coefficient. This matches the intended interpretation of $\rho$ as a measurable “shared-failure” strength: $\rho=0$ corresponds to independent residuals, while $\rho$ near 1 indicates strong groupthink.
Aggregation by averaging.
An internal node with fan-in $b$ receives $\widetilde{X}_{1},\dots,\widetilde{X}_{b}$ and outputs the average
@@
Under the symmetry assumptions above, this is the MSE-optimal linear unbiased aggregator.

### v2

@@ Abstract
Multi-agent systems can improve reliability, yet under a fixed test-time budget they often help, saturate, or even collapse. We develop a deliberately minimal and calibratable theory that predicts these regimes from three binding constraints of modern agent stacks: finite context windows, lossy inter-agent communication, and shared failures among similar agents. Each leaf agent is summarized by a compute–performance scaling exponent $\beta$; communication is captured by a message-length fidelity curve $\gamma(m)$; dependence is captured by an effective shared-error correlation $\rho$; and a context window $W$ imposes hard fan-in limits that make hierarchy necessary. For binary success/failure tasks with majority aggregation, we prove a sharp phase transition for deep $b$-ary trees with correlated inputs and lossy communication: a single scalar $\alpha_{\rho}$ (combining $\gamma(m)$, $\rho$, and fan-in $b$) determines whether weak signal is amplified to a nontrivial fixed point or washed out to chance. In the amplifying regime, we derive an organization exponent $s$ and show that budgeted synergy, i.e., outperforming the best single agent under the same total budget, occurs exactly when $s>\beta$, yielding closed-form compute allocation rules and explicit budget thresholds. We further characterize saturation via a mixing depth and provide a conservative clipped predictor that remains accurate across growth and saturation. A continuous-performance warm-up gives closed-form risks for star, chain, and tree organizations, making correlation- and communication-induced floors explicit and exposing the core design trade-offs in a smooth setting. Finally, we validate the predicted phase boundaries in controlled synthetic simulations and show how the same mechanisms explain the dominant bottlenecks reported in recent large-scale matched-budget studies of LLM agent-system scaling, including context saturation, subcritical error cascades, and diminishing returns at strong baselines.
@@ 1 Introduction
Multi-agent systems are often presented as a reliability primitive: if a single agent is fallible, run many agents, let them interact, and aggregate their outputs. Under a fixed computational budget per task, however, multi-agent coordination is not reliably beneficial. The same additional agents that sometimes improve performance can also lead to saturation or even outright degradation. This paper develops a predictive theory of this brittleness: when multi-agent organization produces budgeted synergy, i.e., outperforming the best single agent under the same total computational budget, and when scaling out must fail.
@@
The LLM era has made this phenomenon impossible to ignore. It is now routine to instantiate multiple LLM-based agents, assign roles, and coordinate them through discussion, critique, or voting in so-called “agent societies” and orchestration frameworks (Brown et al., 2020; Yao et al., 2023; Wu et al., 2023; Li et al., 2023; Chen et al., 2024; Hong et al., 2024). The underlying intuition is classical: ensembling can reduce error (Breiman, 1996; Freund and Schapire, 1997; Dietterich, 2000), and interaction protocols such as debate or amplification aim to turn many weak judgments into a stronger one (Irving et al., 2018; Christiano et al., 2018; Du et al., 2024). At the same time, controlled matched-budget evaluations increasingly reveal strong task- and topology-dependence, including regimes in which multi-agent variants degrade once coordination overhead and information loss dominate (Kim et al., 2025).
@@
Our starting point is that these failures are not implementation accidents but structural consequences of a small number of constraints that repeatedly bind in modern agent stacks. First, agent errors are not independent: multiple instances derived from the same base model and prompt often share systematic failure modes, producing “groupthink” rather than error cancellation. Second, coordination is communication-limited: as systems scale, agents must compress reasoning into bounded-length messages, and the resulting information loss compounds across layers. Third, coordination is context-limited: a central aggregator cannot read arbitrarily many messages under a finite context window, so naive scale-out saturates structurally. Recent work on large-population collaboration and topology learning highlights the same bottleneck: token and context limits are frequently the binding resource (Qian et al., 2025; Zhang et al., 2024).
@@
Today, most agent-system design still proceeds by heuristics: choose a topology, message format, and number of agents, then iterate. In spirit, this resembles the hand-crafted feature engineering era of machine learning: powerful components exist, but system-level behavior depends on ad hoc choices and expensive trial-and-error. Our goal is to push multi-agent design toward an “agentic intelligence from first principles” approach: make the dominant constraints explicit, explain why scaling out works or fails under a fixed computational budget, and derive sharp, quantitative regime boundaries, most notably, a phase transition for budgeted multi-agent synergy.
@@
Figure 1: When does multi-agent organization produce budgeted synergy? An overview of our theoretical framework and core contribution. Left: scaling out multi-agent systems under different topologies (chain, star, and hierarchical tree) subject to a finite context window $W$, which constrains fan-in globally for stars ($Nm\leq W$) and locally for trees ($bm\leq W$). Bottom: three binding constraints of modern agent stacks: finite context windows, lossy communication with fidelity $\gamma(m)$, and shared-failure correlation $\rho$. Right: the competing baseline of scaling up a single agent, governed by a single-agent scaling law with exponent $\beta$. Our theory combines these elements into an effective per-layer gain $\alpha_{\rho}$, which induces an organization exponent $s$ for hierarchical aggregation. Budgeted synergy occurs precisely when $s>\beta$, meaning that scaling out via organization can outperform scaling up a single agent under the same total budget.
@@
Within this framework we analyze three canonical organizations: star (central aggregation), chain (sequential handoff), and hierarchical tree. Our main results characterize when hierarchy is viable and when it is provably counterproductive. For binary success/failure tasks with majority aggregation, we prove a sharp amplification–collapse transition for deep trees: a scalar $\alpha_{\rho}$ (combining communication fidelity, correlation, and fan-in) determines whether weak signal amplifies across layers or is washed out to chance, in the spirit of reconstruction thresholds in noisy broadcasting on trees (Evans et al., 2000; Mossel, 2001; Kesten and Stigum, 1966). In the amplifying regime, we derive an organization exponent $s$ governing growth with the number of leaves and show that budgeted synergy in the growth range occurs exactly when $s>\beta$, yielding closed-form compute allocation and explicit budget thresholds. We also characterize finite-depth saturation via a nontrivial fixed point and mixing-depth guarantees. Finally, we make precise why topology becomes decisive only because constraints bind: absent context limits, central aggregation is information-dominant (data processing inequality) (Cover and Thomas, 2006), and hierarchy is valuable only insofar as it bypasses fan-in limits without crossing into the collapsing regime. Figure 1 illustrates our core contribution.
@@
Our aim in this paper is to theoretically delineate regimes of behavior and identify the mechanisms that govern when multi-agent systems help, saturate, or collapse under realistic constraints. Large-scale matched-budget evaluations with real LLM agent stacks are expensive and can be confounded by prompt engineering and rapidly changing model/tool ecosystems. Accordingly, we validate the sharp regime boundaries via controlled synthetic simulations (Section 6.2) that probe the theory under its own assumptions, and we use external controlled evidence (Kim et al., 2025) to show that the bottlenecks isolated by the theory (context saturation, subcritical collapse, and diminishing returns) are salient in modern deployments.
@@
Section 2 defines the framework, measurable parameters, and the budget and context constraints. Section 3 develops the continuous warm-up. Section 4 presents the binary theory, including the phase transition, small-signal amplification, and saturation; it also compares topologies under context constraints and derives budget thresholds. Section 5 turns the theory into design diagnostics. Section 6 provides empirical touchpoints via external controlled evidence and synthetic sanity checks. We place detailed proofs and optional calibration/evaluation templates in the appendices.
@@ 2.1 Tasks and performance metrics
We study two task primitives that are deliberately simple but theoretically expressive. They are not meant to exhaust the diversity of real-world tasks; rather, they isolate the coordination bottlenecks that arise repeatedly inside more complex multi-agent pipelines. Binary success/failure captures decision-like outcomes (e.g., accept/reject, pass/fail, correctness of a proposed action), while continuous performance on $[0,1]$ captures graded judgments, scores, or calibrated estimates. Together, these two primitives cover a wide range of evaluation signals used in practice and allow clean analysis of aggregation, communication loss, and shared failures. From a theoretical standpoint, any richer task that can be decomposed into local scoring, communication, and aggregation steps must confront the same coordination constraints analyzed here.
The continuous setting serves as a smooth warm-up where many quantities admit closed forms and correlation and communication floors are explicit. The binary setting is not a special case of the continuous one: the nonlinearity of majority vote introduces qualitatively new behavior, including a sharp amplification–collapse phase transition that has no analogue in linear aggregation. Many realistic agent workflows combine both primitives. For example, agents may exchange graded scores or confidence estimates internally, but the final system decision is binary. So analyzing both is necessary to capture the full coordination picture.
@@
$\mu\;:=\;\mathbb{E}[\widehat{Y}Y]\in[-1,1],\qquad\Pr(\widehat{Y}=Y)=\frac{1+\mu}{2}.$ (1)
The bias $\mu$ is a natural summary statistic for weak agents: it measures how far performance is above chance and composes cleanly under simple channel models and aggregation maps. This makes it particularly suitable for analyzing whether small, local improvements are amplified or destroyed by hierarchical organization.
@@
$v\;:=\;\mathbb{E}\big[(\widehat{Y}-Y)^{2}\big],$ (2)
and occasionally report a bounded performance score such as $\mathrm{Perf}=1/(1+v)$, which is monotone in $v$. Continuous metrics capture graded information that is often exchanged within agent systems (scores, confidence levels, value estimates) and lead to linear aggregation dynamics. In this setting, the effect of correlation and communication loss appears as explicit error floors rather than phase transitions, providing intuition that complements the binary analysis.
@@ 2.2 Single-agent capability and scaling
Our central comparison is fixed budget: the same total computational budget $B$ can be allocated either to scale up a single agent (making one agent stronger) or to scale out across many agents (running more agents in parallel). Understanding when scale out can outperform scale up under the same budget is the core question of this paper.
We model per-leaf compute by a single nonnegative knob $x\geq 0$, representing the share of the per-task computational budget allocated to an individual agent. In LLM-based agents, $x$ may correspond to inference tokens devoted to reasoning, the number of samples or self-consistency votes, structured test-time inference procedures (e.g., generate–select or tournament-style algorithms), the number of tool calls, or any other additive per-task resource that trades off directly against parallelism.
We summarize the effect of per-agent compute via a single-agent scaling law. Classical scaling laws are typically reported for training loss as a function of training compute (Kaplan et al., 2020; Hoffmann et al., 2022). Here, however, we use scaling laws in a different but related sense: as an effective description of how a fixed trained agent’s test-time performance improves when allocated additional per-task computational resources. Recent work on test-time or inference-time scaling shows that such improvements can follow systematic power-law or exponential trends in appropriate regimes and metrics, including accuracy, error probability, or calibrated scores (Chen et al., ; Levi, 2025; Snell et al., 2024; Wu et al., 2025). Motivated by these findings, we treat the scaling exponent as a per-task scaling exponent governing scale up, not as a statement about retraining the model.
For binary success/failure tasks, a leaf agent allocated compute $x$ produces a vote $\widehat{Y}$ with bias
@@
where $g(\cdot)$ is increasing and exhibits diminishing returns. For continuous tasks, a leaf produces an estimate $X$ with conditional variance $v_{0}(x)$, where $v_{0}(\cdot)$ decreases with $x$.
@@
with constants $k,c>0$ and a single-agent scaling exponent $\beta>0$. Here “small-signal regime” refers to the operating range of $x$ in which (i) the local power-law approximation is accurate enough for comparison, and (ii) the induced leaf performance is not already saturated. In the binary analysis, this typically corresponds to weak leaf votes (small bias $\mu_{0}(x)$, i.e., accuracy only slightly above chance), which is precisely the regime where the majority update map is well-approximated by its derivative at the origin and where the organization exponent $s$ governs growth. In the continuous warm-up, it corresponds to the pre-saturation range where aggregation has not yet hit the communication/correlation floor. The exponent $\beta$ is not a universal constant: it depends on the model, prompting, tools, test-time procedure, and task family, and is intended to be measured in the operating regime of interest. Our theory does not require the power-law approximation to hold globally or asymptotically; it relies only on the existence of a local scaling exponent $\beta$ in the budget range where the scale-up versus scale-out comparison is made.
@@ 2.3 Communication as a controllable lossy channel
For binary tasks and majority-style protocols, it is natural to treat each child as intending to transmit a single bit of information (its vote or decision), encoded into a message of length $m$. We model the receiver’s decoded bit as passing through a binary symmetric channel (BSC) with reliability $\gamma(m)\in(0,1]$:
$\mu_{\text{recv}}=\gamma(m)\,\mu.$ (5)
Equivalently, $\gamma(m)=1-2\epsilon(m)$, where $\epsilon(m)$ is an effective flip probability.
@@ 2.5 Budget and context constraints
$\text{fan-in}\;\leq\;\Big\lfloor\frac{W}{m}\Big\rfloor.$ (7)
@@ 2.6 Topologies and protocols
Figure 2: From general DAGs to spanning arborescences. A general DAG (a) may contain extra edges (dashed) that add redundancy or dependence, but any feed-forward organization contains a spanning arborescence that routes information from leaves to the output (b).
It is also useful to view more general feed-forward organizations through these motifs. With a finite context window, every node can only aggregate a bounded number of incoming messages, i.e., each local computation is an in-star with fan-in at most $\lfloor W/m\rfloor$. Globally, any connected directed acyclic communication graph contains a spanning arborescence that routes leaf information to the output; star and chain are the two extreme cases (depth $1$ versus fan-in $1$), and bounded-fan-in trees are the minimal structure that scales beyond the star’s context cap. Additional edges or multi-round interaction are valuable mainly when they effectively change the measurable parameters, for example by improving effective fidelity through redundancy or by reducing effective correlation through independent verification, rather than by rerouting the same information along multiple dependent paths. Figure 2 illustrates this reduction from general DAGs to a spanning arborescence.
@@ 3.1 Model assumptions for the continuous setting
This warm-up instantiates the continuous-task primitive from Section 2 using the communication and shared-failure models already defined in Sections 2.3 and 2.4. The only additional assumptions are that leaf estimates are (approximately) unbiased and admit a compute-dependent MSE description.
Let $Y\in[0,1]$ denote the latent target and let each leaf output an estimate $X\in[0,1]$. To keep the warm-up focused on organization rather than estimator bias, we assume conditional unbiasedness,
$\mathbb{E}[X\mid Y]=Y.$ (8)
Write the residual E:=XYE:=X-Y, so $\mathbb{E}[E\mid Y]=0$, and summarize leaf quality by the MSE
$v_{0}(x):=\mathbb{E}\big[(X-Y)^{2}\big]=\mathbb{E}[E^{2}],$ (9)
which depends on the per-leaf compute allocation $x$. In the operating range where we compare scale up versus scale out, we use the scaling approximation
@@
with constants $c>0$ and $\beta>0$ (Section 2.2).
Communication follows the additive distortion model from Section 2.3, with message length $m$; we treat edge distortions as independent across edges and independent of $Y$.
Within any aggregation group, residuals are modeled by the equal-correlation structure from Section 2.4: for children $i\neq j$,
@@
where $\rho\in[0,1)$ is the shared-failure correlation parameter.
Finally, an internal node with fan-in $b$ outputs the average of its received inputs,
@@
Under these symmetry assumptions, averaging is the MSE-optimal linear unbiased aggregator.